OpenStudy (anonymous):

Can someone tell me if I did this correctly? A truck driven in an underground tunnel has mounted on its roof a helium-neon laser that produces a red light who wavelength is 633 nanometers. What is the frequency of the radiation? Here's my work: (I'm posting it)

OpenStudy (anonymous):

\[633\times \frac{ 1 m}{ 1\times10 ^9 nm } = 6.33\times10^{-7}\] meters

OpenStudy (anonymous):

\[6.33\times10^{-7} meters =\frac{ 3.00 x 10^{8} m/s }{ \nu }\]

OpenStudy (anonymous):

Multiply v on both sides so you get rid of it on the right side, then divide the right side by the left side and I get\[.4739\times10^{15} s^{-1}\]

OpenStudy (anonymous):

And that's my answer. Did I do it correctly?

OpenStudy (matt101):

Looks good! You'd probably want to express your answer as 4.74 * 10^14 s^-1

OpenStudy (anonymous):

Okay, thank you! :) I was just worried I made some scientific notation errors along the way.

OpenStudy (matt101):

Nope - the conversion of nm to m is key. In these sorts of questions, you always need to be sure you're working with the same units.